OpenStudy (anonymous):

A car dealership interviewed every driver who came onto their lot on Tuesday. This produced the following results: there were 68 people who owned trucks and 119 people who owned cars. If 17 people owned both a truck and a car and 6 people owned neither, how many people were interviewed?

OpenStudy (anonymous):

when i did it i came up with 176 ppl, is that right

jimthompson5910 (jim_thompson5910):

68 people owned trucks and 17 owned both So 68 - 17 = 51 people owned only trucks ------------ 119 people owned cars and 17 owned both So 119 - 17 = 102 owned only cars So in total, 51+102+17+6 = 176 people were interviewed.

jimthompson5910 (jim_thompson5910):

Yes you got the correct answer. Congrats

OpenStudy (anonymous):

thank you again

jimthompson5910 (jim_thompson5910):

sure thing